OpenStudy (anonymous):

Find the direction angle of v for the following vector: v=-7i-4j (Round to one nearest decimal place as needed.)

OpenStudy (johnweldon1993):

Well it looks like in the 'x' direction (represented by 'i') would be -7...and in the 'y' direction (represented by 'j') would be -4 So \[\large \tan^{-1}(\frac{y}{x})\] \[\large \theta = \tan^{-1}(\frac{-4}{-7})\]

OpenStudy (anonymous):

I got -29.7448813. What do I do next? That's not the right answer.

OpenStudy (johnweldon1993):

why did you get a negative answer? we have a negative / negative so it would be positive

OpenStudy (anonymous):

Oops I meant positive. But I figured it out. Thanks!

OpenStudy (johnweldon1993):

No problem :)